The leakage limit of an aluminum alloy to uniaxial tensile is Y = 320 MPa. The same material also leaks into a complex load described by the tensile state [σij]:

330 95 0 Съ-| 95 70 01 MPa 00150

Consider what leakage criterion (Treska or Mises) best describes the behavior of the material.
Answer #1

Comparing with the matrix

sigmaxx auxy auxz
auyx sigmayy auyz
auzx auzy sigma

We get

auxy=auyx= 95

auxz=auzx= 0

auyz=auzy= 0

sigmaxx= 330

sigmayy= 70

sigmazz= 150

maximum principle stress is given by

sigma=(sigmaxx+sigmayy)/2+(((sigmaxx-sigmayy)/2)2+auxy2)1/2

This gives

sigma1=361

sigma2=180

sigma3=330

and corresponding max shear stresses are :

au1=161

au2=40

au3=90

Now Von mises stress is given by :

sigma^2 =0.5*left ( left ( sigma 1-sigma 2 ight )^2+(sigma2-sigma3)^2 + (sigma 3-sigma1)^2) ight )

therefore sigma= 167.66

Now tresca's criteria states that if the principal stresses are arranged in decreasing order then

Tmar0.5 *

thus Trur = 90.5

Now factor of safety

for von mises is 320/167.66 = 1.91

for Tresca is (320/2)/90.5 = 1.77....... (For isotropic material, Shear stress limit is half that of tensile strength)

Thus Factor of safety for tresca's criteria is less and thus it best describes the behavior of the material.